A person stands in the center of a frictionless fr
Answer #1

m = 20/9.8 kg , M = 420/9.8 kg

From conservation of momentum

MV = (M-m)*vman + mvskate

0 = (420 -20)*vman + 20*15

vman = - 0.75 m/s

velocity of man is 0.75 m/s alongbackward direction